# Morphology of 32 Repeating Fast Radio Burst Sources at Microsecond Time Scales with CHIME/FRB

### v1

@@ 5.3 Multiple FRB Classes?
Kirsten et al. (2024) found that the energy-distribution for the high-energy bursts from FRB 20201124A was consistent with that of non-repeaters, similarly suggesting that non-repeaters might be the brightest burst from a repeater. Most recently, Ould-Boukattine et al. (2024) found similar results for FRB 20220912A. However, Kirsten et al. (2024) did not find any significant morphological or spectral differences between the high-energy bursts from FRB 20201124A and the lower-energy bursts. This could be in-part due to their comparatively smaller bandwidth. We also did not find any distinct morphological differences between the brightest bursts from our sources and the other bursts from the same source. A larger population of bright, repeating sources detected over large bandwidths is needed to explore this in further detail.
@@ 6 Summary and Future Work
In this work, we present for the first time a morphological analysis of the raw voltage data for 118 bursts from 32 unique repeating sources detected by CHIME/FRB. We present spectro-temporal fits for each burst at timescales down to microseconds, along with improved flux and fluence measurements for each burst. We search for correlations amongst the burst fluences, DMs, burst rate, and burst duration. We do not find any significant correlations amongst these parameters, even with improved fluences, burst rates, and burst duration estimates. The lack of a correlation between fluence and extragalactic DM (+MW halo) for our sources suggests that repeating FRBs have intrinsically large luminosity ranges. Additionally, the lack of a correlation between burst rate and total duration may suggest that that burst widths are not correlated with their beaming angle. We also search for an extragalactic DM (+MW halo) and scattering correlation, but again do not find one. A lack of a correlation suggests that the medium responsible for the majority of the extragalactic DM is not the same which primarily contributes to the sources scattering timescales.
@@ acknowledgments
Below, in Figure A1, we include the waterfall plots for a sub-sample of the bursts analyzed in this work. In Table A1, we include the measured burst parameters for each burst. In Figure A3, we show the fluence versus extragalactic DM (+halo) correlation including non-repeating FRBs from Sand et al. (2024). In Figure A4, we show comparisons of the repeating and non-repeating sample, but limit the sample to bursts studied at time resolutions s. In Figure A5 we show the bandwidth versus detection S/N for all sources in our sample for which there are more than five bursts. Lastly, in Table A2, we list updated burst rates for the repeaters in our sample.
@@
| Source | NaaNumber of bursts detected from this source that fall within the FWHM at 600 MHz of the corresponding detection beam. If a source has both a lower and upper transit, only bursts within the upper transit are considered. | ExposurebbTotal exposure to the position of this source. For sources observed twice a day, this corresponds to the exposure solely in the upper transit. | Thresholdccfootnotemark: | Burst Rateeefootnotemark: |
| (hr-1) | (Jy ms) | (hr-1) | ||
| FRB 20201130A | 5 | 43.9 | 16.1 | 0.020 |
| FRB 20191106C | 0 | 2.0 | 30.6 | 0.0+0.12 |
| FRB 20200619A | 4 | 209.0 | 2.3 | 0.061 |
| FRB 20190804E | 4 | 176.6 | 4.4 | 0.027 |
| FRB 20190915D | 4 | 178.5 | 10.7 | 0.007 |
| FRB 20200929C | 0 | 0.0 | 4.6 | – |
| FRB 20201221B | 1 | 1.5 | 16.1 | 0.11 |
| FRB 20200809E | 3 | 914.0 | 10.0 | 0.0012 |
| FRB 20190609C | 2 | 133.7 | 4.4 | 0.018 |
| FRB 20200223B | 4 | 1.6 | 22.3 | 0.27 |
| FRB 20200202A | 2 | 110.7 | 2.1 | 0.07 |
| FRB 20190430C | 3 | 57.7 | 6.4 | 0.036 |
| FRB 20200127B | 0 | 252.5 | 6.7 | 0.0+0.009 |
| FRB 20201114A | 0 | 237.9 | 13.4 | 0.0+0.004 |
| FRB 20190110C | 0 | 88.6 | 1.7 | 0.0+0.2 |
| FRB 20200118D | 1 | 139.1 | 3.8 | 0.011 |
| FRB 20191013D | 1 | 126.4 | 12.4 | 0.0020 |
| FRB 20200926A | 5 | 205.4 | 8.9 | 0.010 |
| FRB 20180910A | 0 | 7400.7 | 10.7 | 0.0+0.0002 |
| FRB 20180814A | 8 | 426.6 | 5.1 | 0.018 |
| FRB 20181030A | 2 | 396.0 | 6.7 | 0.0033 |
| FRB 20181128A | 9 | 184.2 | 6.7 | 0.032 |
| FRB 20181119A | 13 | 266.8 | 1.6 | 0.26 |
| FRB 20190116A | 2 | 127.7 | 9.4 | 0.006 |
| FRB 20190216A | 4 | 322.0 | 6.7 | 0.008 |
| FRB 20190222A | 0 | 307.1 | 0.2 | 0.0+1.55 |
| FRB 20190208A | 15 | 173.8 | 4.5 | 0.102 |
| FRB 20190604A | 3 | 191.1 | 4.6 | 0.018 |
| FRB 20190212A | 0 | 60.8 | 3.4 | 0.0+0.11 |
| FRB 20180908A | – | – | – | – |
| FRB 20190117A | 15 | 127.7 | 20.1 | 0.0145 |
| FRB 20190303A | 1 | 78.5 | 4.1 | 0.017 |
| FRB 20190417A | 12 | 213.5 | 1.6 | 0.32 |
| FRB 20190907A | 5 | 127.5 | 7.0 | 0.024 |
| FRB 20200120E | 7 | 315.6 | 3.8 | 0.033 |
| FRB 20201124A | 23 | 92.8 | 26.1 | 0.021 |
Fluence threshold for the source at the 95% confidence level (only upper transit). ddfootnotetext: Burst rate (in upper transit; if applicable) scaled to a fluence threshold of 5 Jy ms.

### v2

@@ 5.3 Multiple FRB Classes?
Kirsten et al. (2024) found that the energy-distribution for the high-energy bursts from FRB 20201124A was consistent with that of non-repeaters, similarly suggesting that non-repeaters might be the brightest burst from a repeater. Most recently, Pelliciari et al. (2024) and Ould-Boukattine et al. (2024) both found similar results for FRB 20220912A. However, Kirsten et al. (2024) did not find any significant morphological or spectral differences between the high-energy bursts from FRB 20201124A and the lower-energy bursts. This could be in-part due to their comparatively smaller bandwidth. We also did not find any distinct morphological differences between the brightest bursts from our sources and the other bursts from the same source. A larger population of bright, repeating sources detected over large bandwidths is needed to explore this in further detail.
@@ 6 Summary and Future Work
In this work, we present for the first time a morphological analysis of the raw voltage data for 124 bursts from 35 unique repeating sources detected by CHIME/FRB. We present spectro-temporal fits for each burst at timescales down to microseconds, along with improved flux and fluence measurements for each burst. We search for correlations amongst the burst fluences, DMs, burst rate, and burst duration. We do not find any significant correlations amongst these parameters, even with improved fluences, burst rates, and burst duration estimates. The lack of a correlation between fluence and extragalactic DM (+MW halo) for our sources suggests that repeating FRBs have intrinsically large luminosity ranges. Additionally, the lack of a correlation between burst rate and total duration may suggest that that burst widths are not correlated with their beaming angle. We also search for an extragalactic DM (+MW halo) and scattering correlation, but again do not find one. A lack of a correlation suggests that the medium responsible for the majority of the extragalactic DM is not the same which primarily contributes to the sources scattering timescales.
@@ acknowledgments
Below, in Figure A1, we include the waterfall plots for a sub-sample of the bursts analyzed in this work. In Table A1, we include the measured burst parameters for each burst. In Figure A2, we show the maximum burst width from a given source versus its burst rate. In Figure A3, we show the burst duration versus extragalactic DM (+halo) correlation including non-repeating FRBs from Sand et al. (2024). In Figure A4, we show the fluence versus extragalactic DM (+halo) correlation including non-repeating FRBs from Sand et al. (2024). In Figure A5, we show comparisons of the repeating and non-repeating sample, but limit the sample to bursts studied at time resolutions s. In Figure A6 we show the bandwidth versus detection S/N for all sources in our sample for which there are more than five bursts. Lastly, in Table A2, we list updated burst rates for the repeaters in our sample.
@@
Source | NaaNumber of bursts detected from this source that fall within the FWHM at 600 MHz of the corresponding detection beam. If a source has both a lower and upper transit, only bursts within the upper transit are considered. | ExposurebbTotal exposure to the position of this source between 2018 August and 2023 September. For sources observed twice a day, this corresponds to the exposure solely in the upper CHIME transit. | Thresholdccfootnotemark: | Burst RateddBurst rate (in upper transit; if applicable) scaled to a fluence threshold of 5 Jy ms. |
(hr) | (Jy ms) | (hr-1) | ||
FRB 20201130A | 5 | 43.9 | 16.1 | 0.65 |
FRB 20191106C | 0 | 2.0 | 30.6 | 0.0+28.54 |
FRB 20200619A | 4 | 209.0 | 2.3 | 0.0060 |
FRB 20190804E | 4 | 176.6 | 4.4 | 0.019 |
FRB 20190915D | 4 | 178.5 | 10.7 | 0.07 |
FRB 20200929CeeThis source’s primary transit does not fall within the FWHM of any of CHIME’s synthesized beams. Hence, according to the formalism applied in this article, the exposure, number of bursts, and rate for this source is zero. | – | – | – | – |
FRB 20201221B | 1 | 1.5 | 16.1 | 4 |
FRB 20200809E | 3 | 914.0 | 10.0 | 0.009 |
FRB 20190609C | 2 | 133.7 | 4.4 | 0.012 |
FRB 20200223B | 4 | 1.6 | 22.3 | 24 |
FRB 20200202A | 2 | 110.7 | 2.1 | 0.005 |
FRB 20190430C | 3 | 57.7 | 6.4 | 0.07 |
FRB 20200127B | 0 | 252.5 | 6.7 | 0.0+0.02 |
FRB 20201114A | 0 | 237.9 | 13.4 | 0.0+0.07 |
FRB 20190110C | 0 | 88.6 | 1.7 | 0.0+0.009 |
FRB 20200118D | 1 | 139.1 | 3.8 | 0.005 |
FRB 20191013D | 1 | 126.4 | 12.4 | 0.03 |
FRB 20200926A | 5 | 205.4 | 8.9 | 0.058 |
FRB 20180910AffThe particularly high exposure to this source’s position is due to its high (89°) declination. | 0 | 7400.7 | 10.7 | 0.0+0.002 |
FRB 20180814A | 8 | 426.6 | 5.1 | 0.019 |
FRB 20181030A | 2 | 396.0 | 6.7 | 0.008 |
FRB 20181128A | 9 | 184.2 | 6.7 | 0.075 |
FRB 20181119A | 13 | 266.8 | 1.6 | 0.0091 |
FRB 20190116A | 2 | 127.7 | 9.4 | 0.04 |
FRB 20190216A | 4 | 322.0 | 6.7 | 0.019 |
FRB 20190222A | 0 | 307.1 | 0.2 | 0.0+0.0001 |
FRB 20190208A | 15 | 173.8 | 4.5 | 0.073 |
FRB 20190604A | 3 | 191.1 | 4.6 | 0.014 |
FRB 20190212A | 0 | 60.8 | 3.4 | 0.0+0.03 |
FRB 20180908B | 2 | 411.3 | 5.1 | 0.005 |
FRB 20190117A | 15 | 127.7 | 20.1 | 0.95 |
FRB 20190303A | 1 | 78.5 | 4.1 | 0.009 |
FRB 20190417A | 12 | 213.5 | 1.6 | 0.0100 |
FRB 20190907A | 5 | 127.5 | 7.0 | 0.065 |
FRB 20200120E | 7 | 315.6 | 3.8 | 0.015 |
Fluence threshold for the source at the 95% confidence level (only upper transit).